OpenStudy (jmoores):

A spherical fish bowl is half-filled with water. The center of the bowl is C, and the length of segment AB is 12 inches, as shown below. Use Twenty two over for pi. A sphere with diameter 12 inches is drawn. Which of the following can be used to calculate the volume of water inside the fish bowl? (5 points) Select one: a. 1 over 24 over 322 over 7(122)(6) b. 1 over 24 over 322 over 7(123) c. 1 over 24 over 322 over 7(62)(12) d. 1 over 24 over 322 over 7(63)

OpenStudy (jmoores):



OpenStudy (welshfella):

The volume of a sphere = (4/3) pi r^3 here we have a half sphere of water and radius r = 1/2 * 12 = 6 So the volume of water in the bowl = (1/2)(4/3) * (22/7) * 6^3 this should simplify to one of the 4 options.

OpenStudy (jmoores):

Ok, let me try to do it

OpenStudy (jmoores):

I know how to get the numerical answer, but it's asking for something different. Could you explain a little further please? :)

OpenStudy (welshfella):

The options all look a bit weird!!

OpenStudy (jmoores):

Oh I'm so sorry! Let me type them out for you :) Hold on

OpenStudy (welshfella):

are you sure theyt are correct?

OpenStudy (jmoores):

No, I'm not honestly. This is for an online school.

OpenStudy (jmoores):

\[a. 1/2 (4/3)(22/7)(12^2)(6) \]

OpenStudy (jmoores):

\[b. 1/2(4/3)(22/7)(12^3)\]

OpenStudy (jmoores):

\[c. 1/2(4/3)(22/7)(6^2)(12)\]

OpenStudy (jmoores):

\[d. 1/2 (4/3)(22/7)(6^3)\]

OpenStudy (welshfella):

OK so compare them with mine and you'll see which is correct (The * means multiply ) . I didn't have to insert them.

OpenStudy (welshfella):

also 6^3 means 6 to the power 3

OpenStudy (jmoores):

Oh wow. I feel so stupid. I was over complicating it I think

OpenStudy (jmoores):

Thank you so much! Could you help me with a few more?

OpenStudy (welshfella):

Sorry got to go right now..

OpenStudy (jmoores):

Ok, well thank you anyway! :)